Air is a cylindrical is compressed reversible and isothermallay from 95 kPa , 25°C to 290 kPa. The initial volume is 0.162m³. Determine the work input, heat transfer and the entropy Change of the air being compressed

1
Expert's answer
2022-02-25T06:52:02-0500

"P1v1 = p2v2"

"p2\/p1 =r"

"290*10^5\/95*10^5 = r =3.05"

"v2 = v1\/r = 0.162\/3.05"

"= 0.05306m^3"

"W = \int Vdp"

"= V (p2-p1)"

"= 0.053* (290-95)*10^5"

= "1.03mJ"

"R= 5.164*10^3"

"T2 = (290*10^5*0.05306)\/5.164*10^3"

"T2 = 306K" = "33 \degree c"

"\delta S = Cv \log (T2/T1) + R \log(V2/V1)"

= "0.178\log(33/25) + 5.164*10^3(0.05306/.162)"

= 5.763 KJ/Kg/K